Systems Staff Meeting Notes
November 7, 2005
November 7, 2005
Present: Cynthia Bowker, Will Chaney, Suzy Covey, Debra Fetzer, Winston Harris, Bobbie Meng, Tom Minton, Chris Nicolich, Geoffrey Stewart, Ying TangAnnouncements
Review Minutes of 10/31/05
- The hardware crew worked hard over the weekend. Cynthia really gave it her all.
- November 17 Town Meeting - employee recognitions scheduled. All staff are encouraged to attend.
- Bobbie's husband, Carlton, received two more licenses at his job for Putnam County.
- Broken window - in Systems - south window between Ying & Winston's offices. Beer bottle found on outside of window, beer cap found inside. Reported to Facilities, work order generated, PPD is repairing it right now.
Ongoing Systems Updates
- Ann #5 - Tom noted that "Stay" was 'almost as bad' as "Elizabethtown", but "Elizabethtown" is still 'the worse' movie so far this year.
- SW #2 - should be HIPPA, not HEPA.
- Orders - Will has not had time to address those in the queue.
- Meeting occurred last week.
- Public configuration status - Library East swamp, MSL swamp and Maps remain. Overtime scheduled for this Sat-Mon. When the team encounters hardware problems, they will just flag, tag and go. They should not spend time to repair a broken machine.
Software Projects Status
- Backup Exec - still problematic.
- New anti-virus - killed the updated server. Luckily, we had a backup. We switched the servers out. Will is building a WIN 2K3 server to manage McAfee and security policy orchestration. It should push and install updates and remove the old antivirus.
- Greenstone/BBS - still waiting. Won't be happening until Geoffrey returns from vacation.
- Everything is broke: 3 ILLIAD-related problems. The interface between ILL and Business Services is down. ILLIAD needed a restart this morning and the ISO service did not come back up. Ariel is not working very well.
- Chris is working on a project for Access Services - making Electronic Reserves request/submit tool automatic.
- Greenstone Loader - very close to being finished.
- New programmer position - resumes under review.
- Ref Express - transfer of incoming emails -Winston restarted the server and the mail moved to the mailbox, but now they are not moving again.
- There is talk of a communications page for the Library Web site - requested by Dale. Content unknown.
- Trouble Ticket - licenses on SmathersNT2 - online serials catalog - link in Marc record - page not found. The new server is not finding NT2R. Tom will transfer the ticket to Will.
- Meeting was held.